# On the Friendly number of 10

### v1

@@ 4. Proofs of the Main Theorems
Suppose that $n$ has exactly five distinct odd prime factors $p_{1}=5$, $p_{2}$, $p_{3}$, $p_{4}$ and $p_{5}$ with $p_{5}>p_{4}>p_{3}>p_{2}>p_{1}=5$. Then we can write $n=5^{2a_{1}}\cdot p_{2}^{2a_{2}}\cdot p_{3}^{2a_{3}}\cdot p_{4}^{2a_{4}}\cdot p_{5}^{2a_{5}}$ where $a_{1}$, $a_{2}$, $a_{3}$, $a_{4}$, $a_{5}$ are positive integers. Now we will see what are the possible values of $p_{i}$ for $2\leq i\leq 5$. If possible, let $p_{2}>7$ but then by Lemma 2.4 and Lemma 2.5, $I(n)=I(5^{2a_{1}}\cdot p_{2}^{2a_{2}}\cdot p_{3}^{2a_{3}}\cdot p_{4}^{2a_{4}}\cdot p_{5}^{2a_{5}})\leq I(5^{2a_{1}}\cdot 11^{2a_{1}}\cdot p_{3}^{2a_{3}}\cdot p_{4}^{2a_{4}}\cdot p_{5}^{2a_{5}})\leq I(5^{2a_{1}}\cdot 11^{2a_{1}}\cdot 13^{2a_{3}}\cdot 17^{2a_{4}}\cdot 19^{2a_{5}})<\frac{5}{4}\cdot\frac{11}{10}\cdot\frac{13}{12}\cdot\frac{17}{16}\cdot\frac{19}{18}=\frac{46189}{27648}<\frac{9}{5}.$ Hence only possible value of $p_{2}$ is $7$. If possible, let $p_{3}>11$ then by Lemma 2.4 and Lemma 2.5, $I(n)\leq I(5^{2a_{1}}\cdot 7^{2a_{2}}\cdot 13^{2a_{3}}\cdot 17^{2a_{4}}\cdot 19^{2a_{5}})\leq\frac{5}{4}\cdot\frac{7}{6}\cdot\frac{13}{12}\cdot\frac{17}{16}\cdot\frac{19}{18}=\frac{146965}{82944}<\frac{9}{5}$. Hence only possible value of $p_{3}$ is $11$. If possible, let $p_{4}>13$ then again by Lemma 2.4 and Lemma 2.5 $I(n)\leq I(5^{2a_{1}}\cdot 7^{2a_{2}}\cdot 11^{2a_{3}}\cdot 17^{2a_{4}}\cdot 19^{2a_{5}})<\frac{5}{4}\cdot\frac{7}{6}\cdot\frac{11}{10}\cdot\frac{17}{16}\cdot\frac{19}{18}=\frac{24871}{13824}<\frac{9}{5}$. Hence only possible value of $p_{4}$ is $13$. Now if $p_{5}>23$ then by the similar argument we can show that $I(n)<\frac{9}{5}$. Therefore $p_{2}=7$, $p_{3}=11$, $p_{4}=13$ and $p_{5}\in\{17,19,23\}$ are necessary for $I(n)\geq\frac{9}{5}$. Now $p_{5}$ can’t be $17$ by Remark 4.2 as then no $p_{i}$ will be congruent to $1$ modulo $34$. If $p_{5}=23$ then by Corollary 1.3, $p_{3}=11$ is the only prime factor which divides $\sigma(5^{2a_{1}})$ but it allows $71\mid\sigma(5^{2a_{1}})$ and this is absurd as $p_{i}\neq 71$, for $1\leq i\leq 5$. Therefore only possible value of $p_{5}$ is $19$. In this case $p_{3}=11$ and $p_{5}=19$ can divides $\sigma(5^{2a_{1}})$. If $p_{3}=11\mid\sigma(5^{2a_{1}})$ then $71\mid\sigma(5^{2a_{1}})$ but $p_{i}\neq 71$, for $1\leq i\leq 5$. Therefore $p_{5}=19$ must divides $\sigma(5^{2a_{1}})$ but again it allows $31\mid\sigma(5^{2a_{1}})$ which is impossible as $p_{i}\neq 31$, for $1\leq i\leq 5$. Therefore, no prime factor of $n$ can divides $\sigma(5^{2a_{1}})$ that is $\sigma(5^{2a_{1}})$ has prime factors other than $3$, $p_{1}=5$, $p_{2}$, $p_{3}$, $p_{4}$, $p_{5}$. This proves that $n$ can’t be a friend of $10$ and hence $n$ has more than five distinct prime factors.
Case-5:
@@
Suppose that $n$ has exactly six distinct odd prime factors $p_{1}=5$, $p_{2}$, $p_{3}$, $p_{4}$, $p_{5}$ and $p_{6}$ with $p_{6}>p_{5}>p_{4}>p_{3}>p_{2}>p_{1}=5$. Then we can write $n=5^{2a_{1}}\cdot p_{2}^{2a_{2}}\cdot p_{3}^{2a_{3}}\cdot p_{4}^{2a_{4}}\cdot p_{5}^{2a_{5}}\cdot p_{6}^{2a_{6}}$ where $a_{1}$, $a_{2}$, $a_{3}$, $a_{4}$, $a_{5}$, $a_{6}$ are positive integers. Applying similar argument given in Case-4 to get possible values of $p_{i}$, $2\leq i\leq 6$ we obtain the following ,
@@
Subcase-4: For chain(4), $n=5^{2a_{1}}\cdot 7^{2a_{2}}\cdot 11^{2a_{3}}\cdot 17^{2a_{4}}\cdot 19^{2a_{5}}\cdot p_{6}^{2a_{6}}$ where $23\leq p_{6}\leq 2039$ and $a_{1}$, $a_{2}$, $a_{3}$, $a_{4}$, $a_{5}$, $a_{6}$ are positive integers. Since $17\mid n$ by Remark 4.2 there must exists a prime factor of $n$ congruent to $1$ modulo $34$, therefore $p_{6}$ must be congruent to $1$ modulo $34$. Now $5$ can divides $\sigma(11^{2a_{3}})$ and $\sigma(p_{6}^{2a_{6}})$. But if , $5\mid\sigma(11^{2a_{3}})$ then $3221\mid\sigma(11^{2a_{3}})$ that is $p_{6}$ must be $3221$. But $p_{6}$ is not congruent $1$ modulo $34$. Hence 5 must divides $\sigma(p_{6}^{2a_{6}})$. By Proposition 3.1 ($9\mid\sigma(n)$), $3$ must divides one of $\sigma(p_{i}^{2a_{i}})$, for $1\leq i\leq 6$. But $3\nmid\sigma(p_{i}^{2a_{i}})$, for $i=1,3,4$. If $3\mid\sigma(19^{2a_{5}})$ then $127\mid\sigma(19^{2a_{5}})$ that is $p_{6}$ must be $127$ but again it is not congruent $1$ mod $34$. Now it is clear that $3\mid\sigma(7^{2a_{2}})$, we will show that $9$ can’t divides $\sigma(7^{2a_{2}})$. If $9\mid\sigma(7^{2a_{2}})$ then this immediately implies $37\mid\sigma(7^{2a_{2}})$ that is $p_{6}$ must be $37$ but it is not congruent $1$ modulo $34$. Therefore $\sigma(p_{6}^{2a_{6}})$ must be divisible by $3$ as $9\leavevmode\nobreak\ |\leavevmode\nobreak\ \sigma(n)$. Hence $p_{6}$ must be congruent to $1$ modulo $34$ and $\sigma(p_{6}^{2a_{6}})$ must be divisible by $3$ and $5$. Using Theorem 1.5 we get, $p_{6}$ is congruent to $1$ modulo $510$. Therefore $p_{6}$ either $1021$ or $1531$. If $p_{6}=1021$ then $5\mid\sigma(1021^{2a_{6}})$ but this immediately allows $41\mid\sigma(1021^{2a_{6}})$ Which is impossible as $p_{i}\neq 41$, for all $1\leq i\leq 6$. Hence the only choice for $p_{6}$ is $1531$ but again $5\mid\sigma(1531^{2a_{6}})$ implies $691\mid\sigma(1531^{2a_{6}})$ which is impossible by the same argument. Therefore there is no such $p_{6}$ for which $5$ can divides $\sigma(p_{6}^{2a_{6}})$. Hence $n$ can’t be a friend of $10$.

### v2

@@ 4. Proofs of the Main Theorems
Case-4: Suppose that $n$ has exactly five distinct odd prime factors $p_{1}=5$, $p_{2}$, $p_{3}$, $p_{4}$ and $p_{5}$ with $p_{5}>p_{4}>p_{3}>p_{2}>p_{1}=5$. Then we can write $n=5^{2a_{1}}\cdot p_{2}^{2a_{2}}\cdot p_{3}^{2a_{3}}\cdot p_{4}^{2a_{4}}\cdot p_{5}^{2a_{5}}$ where $a_{1}$, $a_{2}$, $a_{3}$, $a_{4}$, $a_{5}$ are positive integers. Now we will see what are the possible values of $p_{i}$ for $2\leq i\leq 5$. If possible, let $p_{2}>7$ but then by Lemma 2.4 and Lemma 2.5, $I(n)=I(5^{2a_{1}}\cdot p_{2}^{2a_{2}}\cdot p_{3}^{2a_{3}}\cdot p_{4}^{2a_{4}}\cdot p_{5}^{2a_{5}})\leq I(5^{2a_{1}}\cdot 11^{2a_{1}}\cdot p_{3}^{2a_{3}}\cdot p_{4}^{2a_{4}}\cdot p_{5}^{2a_{5}})\leq I(5^{2a_{1}}\cdot 11^{2a_{1}}\cdot 13^{2a_{3}}\cdot 17^{2a_{4}}\cdot 19^{2a_{5}})<\frac{5}{4}\cdot\frac{11}{10}\cdot\frac{13}{12}\cdot\frac{17}{16}\cdot\frac{19}{18}=\frac{46189}{27648}<\frac{9}{5}.$ Hence only possible value of $p_{2}$ is $7$. If possible, let $p_{3}>11$ then by Lemma 2.4 and Lemma 2.5, $I(n)\leq I(5^{2a_{1}}\cdot 7^{2a_{2}}\cdot 13^{2a_{3}}\cdot 17^{2a_{4}}\cdot 19^{2a_{5}})\leq\frac{5}{4}\cdot\frac{7}{6}\cdot\frac{13}{12}\cdot\frac{17}{16}\cdot\frac{19}{18}=\frac{146965}{82944}<\frac{9}{5}$. Hence only possible value of $p_{3}$ is $11$. If possible, let $p_{4}>13$ then again by Lemma 2.4 and Lemma 2.5 $I(n)\leq I(5^{2a_{1}}\cdot 7^{2a_{2}}\cdot 11^{2a_{3}}\cdot 17^{2a_{4}}\cdot 19^{2a_{5}})<\frac{5}{4}\cdot\frac{7}{6}\cdot\frac{11}{10}\cdot\frac{17}{16}\cdot\frac{19}{18}=\frac{24871}{13824}<\frac{9}{5}$. Hence only possible value of $p_{4}$ is $13$. Now if $p_{5}>23$ then by the similar argument we can show that $I(n)<\frac{9}{5}$. Therefore $p_{2}=7$, $p_{3}=11$, $p_{4}=13$ and $p_{5}\in\{17,19,23\}$ are necessary for $I(n)\geq\frac{9}{5}$. Now $p_{5}$ can’t be $17$ by Remark 4.2 as then no $p_{i}$ will be congruent to $1$ modulo $34$. If $p_{5}=23$ then by Corollary 1.3, $p_{3}=11$ is the only prime factor which divides $\sigma(5^{2a_{1}})$ but it allows $71\mid\sigma(5^{2a_{1}})$ and this is absurd as $p_{i}\neq 71$, for $1\leq i\leq 5$. Therefore only possible value of $p_{5}$ is $19$. In this case $p_{3}=11$ and $p_{5}=19$ can divides $\sigma(5^{2a_{1}})$. If $p_{3}=11\mid\sigma(5^{2a_{1}})$ then $71\mid\sigma(5^{2a_{1}})$ but $p_{i}\neq 71$, for $1\leq i\leq 5$. Therefore $p_{5}=19$ must divides $\sigma(5^{2a_{1}})$ but again it allows $31\mid\sigma(5^{2a_{1}})$ which is impossible as $p_{i}\neq 31$, for $1\leq i\leq 5$. Therefore, no prime factor of $n$ can divides $\sigma(5^{2a_{1}})$ that is $\sigma(5^{2a_{1}})$ has prime factors other than $3$, $p_{1}=5$, $p_{2}$, $p_{3}$, $p_{4}$, $p_{5}$. This proves that $n$ can’t be a friend of $10$ and hence $n$ has more than five distinct prime factors.
@@
Case-5: Suppose that $n$ has exactly six distinct odd prime factors $p_{1}=5$, $p_{2}$, $p_{3}$, $p_{4}$, $p_{5}$ and $p_{6}$ with $p_{6}>p_{5}>p_{4}>p_{3}>p_{2}>p_{1}=5$. Then we can write $n=5^{2a_{1}}\cdot p_{2}^{2a_{2}}\cdot p_{3}^{2a_{3}}\cdot p_{4}^{2a_{4}}\cdot p_{5}^{2a_{5}}\cdot p_{6}^{2a_{6}}$ where $a_{1}$, $a_{2}$, $a_{3}$, $a_{4}$, $a_{5}$, $a_{6}$ are positive integers. Applying similar argument given in Case-4 to get possible values of $p_{i}$, $2\leq i\leq 6$ we obtain the following ,
@@
Subcase-4: For chain(4), $n=5^{2a_{1}}\cdot 7^{2a_{2}}\cdot 11^{2a_{3}}\cdot 17^{2a_{4}}\cdot 19^{2a_{5}}\cdot p_{6}^{2a_{6}}$ where $23\leq p_{6}\leq 2039$ and $a_{1}$, $a_{2}$, $a_{3}$, $a_{4}$, $a_{5}$, $a_{6}$ are positive integers. Since $17\mid n$ by Remark 4.2 there must exists a prime factor of $n$ congruent to $1$ modulo $34$, therefore $p_{6}$ must be congruent to $1$ modulo $34$. Now $5$ can divides $\sigma(11^{2a_{3}})$ and $\sigma(p_{6}^{2a_{6}})$. But if , $5\mid\sigma(11^{2a_{3}})$ then $3221\mid\sigma(11^{2a_{3}})$ that is $p_{6}$ must be $3221$. But $p_{6}$ is not congruent $1$ modulo $34$. Hence 5 must divides $\sigma(p_{6}^{2a_{6}})$. By Proposition 3.1 ($9\mid\sigma(n)$), $3$ must divides one of $\sigma(p_{i}^{2a_{i}})$, for $1\leq i\leq 6$. But $3\nmid\sigma(p_{i}^{2a_{i}})$, for $i=1,3,4$. If $3\mid\sigma(19^{2a_{5}})$ then $127\mid\sigma(19^{2a_{5}})$ that is $p_{6}$ must be $127$ but again it is not congruent $1$ mod $34$. Now it is clear that $3\mid\sigma(7^{2a_{2}})$, we will show that $9$ can’t divides $\sigma(7^{2a_{2}})$. If $9\mid\sigma(7^{2a_{2}})$ then this immediately implies $37\mid\sigma(7^{2a_{2}})$ that is $p_{6}$ must be $37$ but it is not congruent $1$ modulo $34$. Therefore $\sigma(p_{6}^{2a_{6}})$ must be divisible by $3$ as $9~{}|~{}\sigma(n)$. Hence $p_{6}$ must be congruent to $1$ modulo $34$ and $\sigma(p_{6}^{2a_{6}})$ must be divisible by $3$ and $5$. Using Theorem 1.5 we get, $p_{6}$ is congruent to $1$ modulo $510$. Therefore $p_{6}$ either $1021$ or $1531$. If $p_{6}=1021$ then $5\mid\sigma(1021^{2a_{6}})$ but this immediately allows $41\mid\sigma(1021^{2a_{6}})$ Which is impossible as $p_{i}\neq 41$, for all $1\leq i\leq 6$. Hence the only choice for $p_{6}$ is $1531$ but again $5\mid\sigma(1531^{2a_{6}})$ implies $691\mid\sigma(1531^{2a_{6}})$ which is impossible by the same argument. Therefore there is no such $p_{6}$ for which $5$ can divides $\sigma(p_{6}^{2a_{6}})$. Hence $n$ can’t be a friend of $10$.